Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, January 06, 1999

The following is a brief summary of the environmental monitoring activities that were completed on January 06, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/06/99 (Wednesday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 6 January 1999: Backfill operations continued along Peace Valley Road, near the entrance to the Flying J. Crews at this location also worked on welding 8" pipe to transfer hydrotest water from one section of pipe to the next. Fiber optic crews pulled cable from just north of the Flying J to south of that point. Coating moved forward with work at the Warbler Skip, with paving crews working on repairing Lebec Road. A sizing plate was successfully run between the I-5 Tejon bore and the Warbler Skip. Backfill of the Tejon water pipe and PPSI FOC began today. Crews continued to work on clean up along the bench. A joint was welded, x-rayed, and coated above ground on upper Grapevine Road. The joint was lowered into the ditch and tied in to the lower mainline. On the lower portion of that road, crews continued to prepare the road for more paving. Work at Grapevine Creek consisted of more earthmoving and stringing of pipe nearby. Surplus pipe was strung, welded, and x-rayed along Highway 166 again today. Minor clean up began along the I-5 frontage road. Work occurred at both the Grapevine and Emidio Stations.

Segment 2B (Castaic to Sylmar). Report for 01-06-99. Valve and electrical work continued at the vault locations on this segment. Asphalt grinding, removal and repaving took place over the trench in a segment of The Old Road immediately south of Weldon Canyon Road. Non-compliance report #220040 was issued regarding construction activities continuing in the roadway, on the Caltrans ROW south of Weldon Canyon Road, beyond permitted hours.

Segment 3 (Sylmar to Los Angeles). Report for 01-06-99. Tie-in activities (welding and NDT) took place at Spring St (cornfield), and electrical work occurred at that location. Slurry trucks filled electrical ditch on the south side of the street here. Electrical work took place at the vault site at Balboa. Electrical work took place at the intersection of Highland Avenue and San Fernando Road and street saws were used in the intersection. More electrical work occurred near the Astoria intersection. Tie-in activity continued at the trench just south of Balboa Ave.

Segment 5 (Lynwood to Long Beach): Report for 1-6-99. Crews set the bore equipment into the pit and backfilled around the bore pit at the Edison station site. Near Tosco on Alameda St. crews coated the 16" pipe then backfilled the trenchline. Within the station, crews continued electrical work and x-rayed the drain lines for the skids. Segment 4B (Lynwood to El Segundo): Electrical crews continued with final checks and red concrete was poured over underground electrical lines at the El Segundo Station. Likewise, crews at Lynwood station progressed with station lighting and brick work. Further, crews were on site to estimate materials required to coat the exposed above ground pipe.

Inspection Summar

In the past 24 hours, 4 inspection event(s) were reported for 4 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

Spread 2: NCR Report(s) Issued

NCR# 1990114 Monitor's Name: Chris Fazzari Inspection Date: 1/06/99  
Inspection Code: 1 Castaic to Sylmar
Begin Station: 33 End Station: 64 Landowner: County Activity: General

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page